# Excitonic Theory of the Ultrafast Optical Response of 2D-Quantum-Confined Semiconductors at Elevated Densities

### v1

@@ III Excitonic Equations of Motion
Here, $E_{\mu}^{\xi}=E_{\mu,\mathbf{Q}=\mathbf{0}}^{\xi}$ is the excitonic energy with excitonic quantum number $\mu$ at zero center-of-mass momentum $\mathbf{Q}$ and spin-valley index $\xi$, $\hbar\omega_{\text{P}}$ is the center frequency of the exciting optical pulse in a rotating frame, $\gamma_{\text{nrad}}$ is the total non-radiative homogeneous dephasing $\gamma_{\text{nrad}}=\gamma_{\text{X-phon}}+\gamma_{\text{res}}$ by exciton-phonon interaction $\gamma_{\text{X-phon}}$ Selig et al. (2016); Lengers et al. (2020), cf. circled arrow denoted by $\Gamma$” in Fig. 4, and by other residual contributions $\gamma_{\text{res}}$ via, e.g., scattering with uniformly distributed impurities Ma and Jena (2014); Jena and Konar (2007). Inhomogeneous contributions are not considered, as their evaluation entails an ensemble average over many spatially dependent and randomly fluctuating slab- Glutsch and Bechstedt (1994); Thränhardt et al. (2003) or substrate thicknesses Raja et al. (2019) or non-uniformly distributed impurities, which greatly increases the numerical complexity. $\hbar\Omega^{cv}=\left(\hbar\Omega^{vc}\right)^{*}$ is the Rabi energy:
@@
with transition dipole moment $\mathbf{d}^{cv,\xi}$ Aversa and Sipe (1995); Pfalz et al. (2005); Mkrtchian et al. (2019) and optical field in circular basis in rotating wave approximation:
@@
and neglect any contributions due to reradiation Knorr et al. (1996).
Throughout the manuscript, we assume normal incidence with $\mathbf{E}_{\mathbf{Q}}=\mathbf{E}\delta_{\mathbf{Q},\mathbf{0}}$, i.e., we always assume spatial in-plane homogeneity and neglect any diffusion Hess and Kuhn (1996a, b), as such effect is negligible during the ultrafast timescales considered in this manuscript Wagner et al. (2021); Rosati et al. (2021); Zipfel et al. (2020); Zhao et al. (2003).
In Fig. 4, we depict the action of the optical pulse ($\Omega$) annihilating an excitonic transition $P$ in first order and the coupling to $B$ and $N$ in second order: Here, the $y$-axis denotes the total annihilated electron-hole pairs described by the corresponding correlation $P$, $B$, NN etc. and the $x$-axis denotes the total electron-hole pair number $n$, cf. also Eq. (11). Note again, that – in contrast to, e.g., Refs. Victor et al. (1995); Axt and Stahl (1994); Axt et al. (1996a) – we always use the correlated excitonic operator expectation values throughout this work.
@@
where the first line denotes the free excitonic contribution. We note, that, on this level, exciton-phonon interaction (second line in Eq. (27)) provides an additional important contribution, which, in second-order Born-Markov approximation, can be written as Axt et al. (1996b); Thränhardt et al. (2000); Selig et al. (2018):
@@
Here, the first line denotes the phonon-assisted formation of excitonic occupations/intraexcitonic coherences, process $\Gamma P^{*}$” in Fig. 4, while the remaining terms denote phonon-assisted scattering and thermalization, process $\Gamma$” denoted by a circled arrow. The phonon-assisted scattering rates $\Gamma^{\text{in/out}}$ can be deduced from, e.g., Refs. Selig et al. (2018); Thränhardt et al. (2000). Since we focus on light-matter- and Coulomb interaction at elevated densities during intense optical fields beyond the $\chi^{(3)}$-limit, we neglect incoherent exciton-phonon scattering in the following. Nevertheless, we want to emphasize, that exciton-phonon scattering can become important in the incoherent limit at nonlinear exciton densities. For a detailed analysis of the occurring fermionic, bosonic and exchange contributions we refer to Refs. Katzer et al. (2023, 2024). Direct optical-field pump-contributions to the excitonic occupations/intraexcitonic coherences, denoted by $\Omega$”, vanish: $\mathrm{i}\hbar\partial_{t}N\vphantom{N}_{\sigma,\lambda,\mathbf{Q}}^{\xi,\xi^{\prime}}\Big|_{\mathbf{E}^{2},\Omega}=0$. Note, that if we additionally considered intraband light-matter interaction Huttner et al. (2017), intraexcitonic coherences and excitonic drift can occur already in second-order via, e.g., terahertz fields Kira et al. (2001); Stroucken et al. (2021); Kira et al. (2003).
The equations of motion for the two-exciton transitions read Katsch et al. (2019, 2020); Schumacher et al. (2006):
@@
Here, $E_{B,\pm,\zeta}^{\xi,\xi^{\prime}}$ is the two-exciton dispersion with two-exciton quantum number $\zeta$, configuration $\pm$ and two-exciton dephasing $\gamma_{B}$, which we set as $\gamma_{B}=2\gamma_{\text{nrad}}$. Note, that the dephasing of bound biexcitons can differ from this assumption Langbein and Hvam (2000). $\Phi^{*}{\vphantom{\varphi}}_{B,\pm,\zeta,\sigma,\eta,\mathbf{K}}^{\text{L},\xi,\xi,\xi^{\prime},\xi^{\prime}}$ is the two-exciton wave function as a solution of the two-exciton Schrödinger equation in Eq. (S18), $S$ is the projection matrix in Eq. (S23) and $V_{1}$ and $V_{2}$ are exciton-exciton matrix elements, cf. Eq. (S24) and Eq. (S25). The ”++” configuration, which can occur within equal valleys $\xi=\xi^{\prime}$ or distinct valleys $\xi\neq\xi^{\prime}$, hosts only continuum states. In contrast, the ”-” configuration, where only two excitonic transitions from different spin-valleys $\xi\neq\xi^{\prime}$ are involved, as the contribution within equal valleys $\xi=\xi^{\prime}$ vanishes, hosts one bound state, the bound biexciton, and a continuum. In general, two-exciton transitions are induced via a product of two excitonic transitions $P$ via the exciton-exciton scattering encoded in the matrix elements $V_{1/2}$. The corresponding process is depicted in Fig. 4 via ”VPVP”.
@@
where the first two lines encode incoherent band-gap reduction and excitonic phase-space filling, cf. process ”VPVP” denoted by an arrow from $N$ to $P$ in Fig. 5, mediated via the Coulomb matrix element $U_{2/3/4}$, cf. Eqs. (S39)–(S41), the third line describes incoherent Pauli-blocking, cf. process $\Omega$” in Fig. 5, via the matrix elements $D_{1/2}$, cf. Eq. (S28) and Eq. (S29).
@@
where the first two lines describe the interaction with two-exciton transitions $B\vphantom{B}$, cf. processes $VP^{*}$” in Fig. 5, and the last two lines describe the interaction with exciton-two-exciton transitions $Z\vphantom{Z}$, cf. processes ”VV” with the exciton-exciton matrix elements $V_{1/2}$, cf. Eq. (S24) and Eq. (S25).
@@
Here, $E_{R,\pm,\zeta,\rho,\mathbf{Q}}^{\xi_{1},\xi_{2},\xi_{3},\xi_{4},\xi_{5},\xi_{6}}$ is the exciton-two-exciton dispersion for configuration $\pm$, exciton-two-exciton quantum number $\zeta$, momentum $\mathbf{Q}$ and $\gamma_{Z}$ is the exciton-biexciton dephasing, which we set as $\gamma_{Z}=3\gamma_{\text{nrad}}$. Note, that the dephasing of the bound exciton-biexciton may differ from this assumption Langbein and Hvam (2000). Similar to Eq. (18), the ”++” configuration, which exists for intra- and intervalley configurations, hosts exciton-two-exciton continuum states, while the ”-” configuration, which vanishes for intravalley interactions, hosts one bound state and a continuum. The main difference to the two-exciton transitions in Eq. (18) is, that the exciton-two-exciton transitions exhibit a partly incoherent nature, as they are induced by a product of intraexcitonic coherences/excitonic occupations $N_{\mathbf{Q}}$ and excitonic transitions $P$ mediated via the exciton-exciton matrix elements $V_{1/3/4}$. The corresponding process ”VNVN” is denoted by an arrow from $P$ to $Z$. We note, that ”VPVP”, denoted by an arrow from $N$ to $Z$ in Fig. 5 denotes the same process, but from a different viewpoint. For that matter, the exciton-two-exciton wave functions $\Phi^{*}{\vphantom{\varphi}}_{R,\pm,\zeta;\mu,\delta,\mathbf{K}}^{\text{L},\xi_{3},\xi_{4},\xi_{5},\xi_{6}}\left(\mathbf{Q}\right)$ solving the exciton-two-exciton Schrödinger equation in Eq. (S19) acquire an additional dependence on the center-of-mass momentum $\mathbf{Q}$ corresponding to the exciton-two-exciton-inducing intraexcitonic coherence/incoherent excitonic occupation $N_{\mathbf{Q}}{}$. The exciton-exciton matrix elements $V_{1/3/4}$ entering Eq. (33) are given in Eq. (S24), Eq. (S26) and Eq. (S27), respectively.
Compared to second order, incoherent occupations/intraexcitonic coherences and two-exciton transitions couple back to the excitonic transitions via many-body optical and/or Coulomb interaction processes, cf. Fig. 5. Additionally, the partially incoherent exciton-two-exciton transitions, cf. Eq. (20), which emerge in third order, cf. Eq. (33), and, at the same time, couple back to the excitonic transitions, cf. Eq. (32), serve as an important bridge between the coherent and incoherent dynamics. In particular, in two-pulse pump-probe experiments, they adopt the role of the fully coherent two-exciton transitions $B$ in Eq. (18), when all pump-induced coherent excitonic transitions $P$ have been decayed and the system is in a fully incoherent state $N_{\mathbf{Q}}{}$. Hence, if we combine the equations of motion up to third order with exciton-phonon scattering (and density-independent Coulomb scattering such as intervalley exchange Selig et al. (2019, 2020); Yu and Wu (2014); Qiu et al. (2015) or Dexter interaction Dogadov et al. (2025); Berghäuser et al. (2018)), a comprehensive description of pump-probe experiments from the coherent to the incoherent regime valid for small to moderate pump fluences can be obtained.
@@
Beyond a fully coherent limit, Coulomb interaction causes a coupling between intraexcitonic coherences and excitonic occupations, cf. process ”VNVN” in Fig. 6:
@@
Therefore, there does not exist a direct coupling between excitonic occupations, but excitonic occupations at quantum number $\mu$ can induce intraexcitonic coherences at quantum number $\mu,\nu$, which in turn couple back to excitonic occupations at quantum number $\nu$ creating potential excitonic oscillations. In Refs. Wang et al. (2007); Wang and Dignam (2009), similar expressions have been derived in the limit of vanishing center-of-mass momentum, where the interaction of excitonic occupations and intraexcitonic coherences involving higher-lying excitonic states has been shown to cause an oscillatory behavior in the $1s$ excitonic occupation in time.
@@
Eq. (43) and Eq. (44) exhibit very lengthy expressions, so that we refer to Eq. (S66) and Eq. (S67) in the SM for their explicit forms. Again, similar to Eq. (34), Eq. (43) and Eq. (44) induce incoherent, optically dark excitonic occupations out of the light cone due to optical and Coulomb interaction processes via intermediate two-exciton transitions or exciton-two-exciton transitions. Similar contributions to the coherent formation of excitonic occupations outside the light cone at $\mathbf{Q}>\mathbf{0}$ have been found in Ref. Yang et al. (2004), where a fully bosonic excitonic description has been applied. In particular, the fully coherent contributions in Eq. (44) (first terms $\sim P^{*}P^{*}B$) resemble the terms found in Ref. Yang et al. (2004). However, in our numerical simulations, it turns out, that the most important contributions to the coherent excitation of excitonic occupations out of the light cone are the fermionic Pauli-blocking contributions in Eq. (34) and Eq. (35). Eq. (44) constitute only minor corrections, which we found to attenuate rather than increase the overall coherently excited excitonic occupation out of the light cone. This observation is in line with the argumentation in Ref. Axt et al. (2001a), where similar expressions in a Wannier basis have been derived.
@@
Moreover, in a full treatment without a Markovian elimination, the additionally appearing fourth-order terms in the equations of motion of $B$ give rise to optical Pauli-blocking and Coulomb energy renormalizations of the biexcitons and a coupling between biexcitons $B$ and biexcitonic occupations, if correlated eight-operator expectation values are explicitly considered as well. These contributions could be necessary to explain power-dependent biexcitonic photoluminescence signatures Ye et al. (2018) opening up the possibility of a theoretical understanding of the biexciton gain Hayamizu et al. (2007). However, their evaluation is beyond the scope of this manuscript.
We also note, that incoherent exciton scattering processes such as Auger scattering Steinhoff et al. (2021); Erkensten et al. (2021) and density-dependent exciton-phonon scattering Katzer et al. (2024, 2023) are also fourth-order processes, which can be straightforwardly included in our theory by adding the respective interaction Hamiltonians in Eq. (1).

### v2

@@ III Excitonic Equations of Motion
Here, $E_{\mu}^{\xi}=E_{\mu,\mathbf{Q}=\mathbf{0}}^{\xi}$ is the excitonic energy with excitonic quantum number $\mu$ at zero center-of-mass momentum $\mathbf{Q}$ and spin-valley index $\xi$, $\hbar\omega_{\text{P}}$ is the center frequency of the exciting optical pulse in a rotating frame, $\gamma_{\text{nrad}}$ is the total non-radiative homogeneous dephasing $\gamma_{\text{nrad}}=\gamma_{\text{X-phon}}+\gamma_{\text{res}}$ by exciton-phonon interaction $\gamma_{\text{X-phon}}$ [109, 79], cf. circled arrow denoted by $\Gamma$” in Fig. 4, and by other residual contributions $\gamma_{\text{res}}$ via, e.g., scattering with uniformly distributed impurities [83, 56]. Inhomogeneous contributions are not considered, as their evaluation entails an ensemble average over many spatially dependent and randomly fluctuating slab- [45, 123] or substrate thicknesses [98] or non-uniformly distributed impurities, which greatly increases the numerical complexity. $\hbar\Omega^{cv}=\left(\hbar\Omega^{vc}\right)^{*}$ is the Rabi energy:
@@
with transition dipole moment $\mathbf{d}^{cv,\xi}$ [4, 94, 88] and optical field in circular basis in rotating wave approximation:
@@
and neglect any contributions due to reradiation [67].
Throughout the manuscript, we assume normal incidence with $\mathbf{E}_{\mathbf{Q}}=\mathbf{E}\delta_{\mathbf{Q},\mathbf{0}}$, i.e., we always assume spatial in-plane homogeneity and neglect any diffusion [49, 50], as such effect is negligible during the ultrafast timescales considered in this manuscript [128, 101, 137, 136].
In Fig. 4, we depict the action of the optical pulse ($\Omega$) annihilating an excitonic transition $P$ in first order and the coupling to $B$ and $N$ in second order: Here, the $y$-axis denotes the total annihilated electron-hole pairs described by the corresponding correlation $P$, $B$, NN etc. and the $x$-axis denotes the total electron-hole pair number $n$, cf. also Eq. (11). Note again, that – in contrast to, e.g., Refs. [127, 6, 7] – we always use the correlated excitonic operator expectation values throughout this work.
@@
where the first line denotes the free excitonic contribution. We note, that, on this level, exciton-phonon interaction (second line in Eq. (28)) provides an additional important contribution, which, in second-order Born-Markov approximation, can be written as [10, 124, 110]:
@@
Here, the first line denotes the phonon-assisted formation of excitonic occupations/intraexcitonic coherences, process $\Gamma P^{*}$” in Fig. 4, while the remaining terms denote phonon-assisted scattering and thermalization, process $\Gamma$” denoted by a circled arrow. The phonon-assisted scattering rates $\Gamma^{\text{in/out}}$ can be deduced from, e.g., Refs. [110, 124]. Since we focus on light-matter- and Coulomb interaction at elevated densities during intense optical fields beyond the $\chi^{(3)}$-limit, we neglect incoherent exciton-phonon scattering in the following. Nevertheless, we want to emphasize, that exciton-phonon scattering can introduce important corrections in the incoherent limit at nonlinear exciton densities. For a detailed analysis of the occurring fermionic, bosonic and exchange contributions we refer to Refs. [62, 61]. Direct optical-field pump-contributions to the excitonic occupations/intraexcitonic coherences, denoted by $\Omega$”, vanish: $\mathrm{i}\hbar\partial_{t}N\vphantom{N}_{\sigma,\lambda,\mathbf{Q}}^{\xi,\xi^{\prime}}\Big|_{\mathbf{E}^{2},\Omega}=0$. Note, that if we additionally considered intraband light-matter interaction [52], intraexcitonic coherences and excitonic drift can occur already in second-order via, e.g., terahertz fields [63, 118, 64].
The equations of motion for the two-exciton transitions read [59, 60, 108]:
@@
Here, $E_{B,\pm,\zeta}^{\xi,\xi^{\prime}}$ is the two-exciton dispersion with two-exciton quantum number $\zeta$, configuration $\pm$ and two-exciton dephasing $\gamma_{B}$, which we set as $\gamma_{B}=2\gamma_{\text{nrad}}$. Note, that the dephasing of bound biexcitons can differ from this assumption [75]. $\Phi^{*}{\vphantom{\varphi}}_{B,\pm,\zeta,\sigma,\eta,\mathbf{K}}^{\text{L},\xi,\xi,\xi^{\prime},\xi^{\prime}}$ is the two-exciton wave function as a solution of the two-exciton Schrödinger equation in Eq. (S18), $S$ is the projection matrix in Eq. (S23) and $V_{1}$ and $V_{2}$ are exciton-exciton matrix elements, cf. Eq. (S24) and Eq. (S25). The “++” configuration, which can occur within equal valleys $\xi=\xi^{\prime}$ or distinct valleys $\xi\neq\xi^{\prime}$, hosts only continuum states. In contrast, the “-” configuration, where only two excitonic transitions from different spin-valleys $\xi\neq\xi^{\prime}$ are involved, as the contribution within equal valleys $\xi=\xi^{\prime}$ vanishes, hosts one bound state, the bound biexciton, and a continuum. In general, two-exciton transitions are induced via a product of two excitonic transitions $P$ via the exciton-exciton scattering encoded in the matrix elements $V_{1/2}$. The corresponding process is depicted in Fig. 4 via “VPVP”.
@@
where the first two lines encode incoherent band-gap reduction and excitonic phase-space filling, cf. process “VPVP” denoted by an arrow from $N$ to $P$ in Fig. 5, mediated via the Coulomb matrix element $U_{2/3/4}$, cf. Eqs. (S39)–(S41), the third line describes incoherent Pauli-blocking, cf. process $\Omega$” in Fig. 5, via the matrix elements $D_{1/2}$, cf. Eq. (S28) and Eq. (S29).
@@
where the first two lines describe the interaction with two-exciton transitions $B\vphantom{B}$, cf. processes $VP^{*}$” in Fig. 5, and the last two lines describe the interaction with exciton-two-exciton transitions $Z\vphantom{Z}$, cf. processes “VV” with the exciton-exciton matrix elements $V_{1/2}$, cf. Eq. (S24) and Eq. (S25).
@@
Here, $E_{R,\pm,\zeta,\rho,\mathbf{Q}}^{\xi_{1},\xi_{2},\xi_{3},\xi_{4},\xi_{5},\xi_{6}}$ is the exciton-two-exciton dispersion for configuration $\pm$, exciton-two-exciton quantum number $\zeta$, momentum $\mathbf{Q}$ and $\gamma_{Z}$ is the exciton-biexciton dephasing, which we set as $\gamma_{Z}=3\gamma_{\text{nrad}}$. Note, that the dephasing of the bound exciton-biexciton may differ from this assumption [75]. Similar to Eq. (19), the “++” configuration, which exists for intra- and intervalley configurations, hosts exciton-two-exciton continuum states, while the “-” configuration, which vanishes for intravalley interactions, hosts one bound state and a continuum. The main difference to the two-exciton transitions in Eq. (19) is, that the exciton-two-exciton transitions exhibit a partly incoherent nature, as they are induced by a product of intraexcitonic coherences/excitonic occupations $N_{\mathbf{Q}}$ and excitonic transitions $P$ mediated via the exciton-exciton matrix elements $V_{1/3/4}$. The corresponding process “VNVN” is denoted by an arrow from $P$ to $Z$. We note, that “VPVP”, denoted by an arrow from $N$ to $Z$ in Fig. 5 denotes the same process, but from a different viewpoint. For that matter, the exciton-two-exciton wave functions $\Phi^{*}{\vphantom{\varphi}}_{R,\pm,\zeta;\mu,\delta,\mathbf{K}}^{\text{L},\xi_{3},\xi_{4},\xi_{5},\xi_{6}}\left(\mathbf{Q}\right)$ solving the exciton-two-exciton Schrödinger equation in Eq. (S19) acquire an additional dependence on the center-of-mass momentum $\mathbf{Q}$ corresponding to the exciton-two-exciton-inducing intraexcitonic coherence/incoherent excitonic occupation $N_{\mathbf{Q}}{}$. The exciton-exciton matrix elements $V_{1/3/4}$ entering Eq. (34) are given in Eq. (S24), Eq. (S26) and Eq. (S27), respectively.
Compared to second order, incoherent occupations/intraexcitonic coherences and two-exciton transitions couple back to the excitonic transitions via many-body optical and/or Coulomb interaction processes, cf. Fig. 5. Additionally, the partially incoherent exciton-two-exciton transitions, cf. Eq. (21), which emerge in third order, cf. Eq. (34), and, at the same time, couple back to the excitonic transitions, cf. Eq. (33), serve as an important bridge between the coherent and incoherent dynamics. In particular, in two-pulse pump-probe experiments, they adopt the role of the fully coherent two-exciton transitions $B$ in Eq. (19), when all pump-induced coherent excitonic transitions $P$ have been decayed and the system is in a fully incoherent state $N_{\mathbf{Q}}{}$. Hence, if we combine the equations of motion up to third order with exciton-phonon scattering (and density-independent Coulomb scattering such as intervalley exchange [112, 111, 135, 96] or Dexter interaction [35, 15]), a comprehensive description of pump-probe experiments from the coherent to the incoherent regime valid for small to moderate pump fluences can be obtained.
@@
Beyond a fully coherent limit, Coulomb interaction causes a coupling between intraexcitonic coherences and excitonic occupations, cf. process “VNVN” in Fig. 6:
@@
Therefore, there does not exist a direct coupling between excitonic occupations, but excitonic occupations at quantum number $\mu$ can induce intraexcitonic coherences at quantum number $\mu,\nu$, which in turn couple back to excitonic occupations at quantum number $\nu$ creating potential excitonic oscillations. In Refs. [130, 129], similar expressions have been derived in the limit of vanishing center-of-mass momentum, where the interaction of excitonic occupations and intraexcitonic coherences involving higher-lying excitonic states has been shown to cause an oscillatory behavior in the $1s$ excitonic occupation in time.
@@
Eq. (44) and Eq. (45) exhibit very lengthy expressions, so that we refer to Eq. (S66) and Eq. (S67) in the SM for their explicit forms. Again, similar to Eq. (35), Eq. (44) and Eq. (45) induce incoherent, optically dark excitonic occupations out of the light cone due to optical and Coulomb interaction processes via intermediate two-exciton transitions or exciton-two-exciton transitions. Similar contributions to the coherent formation of excitonic occupations outside the light cone at $\mathbf{Q}>\mathbf{0}$ have been found in Ref. [133], where a fully bosonic excitonic description has been applied. In particular, the fully coherent contributions in Eq. (45) (first terms $\sim P^{*}P^{*}B$) resemble the terms found in Ref. [133]. However, in our numerical simulations, it turns out, that the most important contributions to the coherent excitation of excitonic occupations out of the light cone are the fermionic Pauli-blocking contributions in Eq. (35) and Eq. (36). Eq. (45) constitute only minor corrections, which we found to attenuate rather than increase the overall coherently excited excitonic occupation out of the light cone. This observation is in line with the argumentation in Ref. [8], where similar expressions in a Wannier basis have been derived.
@@
Moreover, in a full treatment without a Markovian elimination, the additionally appearing fourth-order terms in the equations of motion of $B$ give rise to optical Pauli-blocking and Coulomb energy renormalizations of the biexcitons and a coupling between biexcitons $B$ and biexcitonic occupations, if correlated eight-operator expectation values are explicitly considered as well. These contributions could be necessary to explain power-dependent biexcitonic photoluminescence signatures [134] opening up the possibility of a theoretical understanding of the biexciton gain [48]. However, their evaluation is beyond the scope of this manuscript.
We also note, that incoherent exciton scattering processes such as Auger scattering [116, 39] and density-dependent exciton-phonon scattering [61, 62] are also fourth-order processes, which can be straightforwardly included in our theory by adding the respective interaction Hamiltonians in Eq. (1).